# Solving the Granularity Mismatch: Hierarchical Preference Learning for Long-Horizon LLM Agents

### v1

@@ 4.4 Ablation on Hierarchical DPO Losses (RQ4)
To investigate the individual contribution of each component in our hierarchical framework, we conduct an ablation study on the three DPO losses, with results shown in Figure 5. The results reveal that while all three loss components, trajectory, action, and group, contribute positively to the final performance, the group-level DPO is demonstrably the most critical. Across all benchmarks, removing the group-level signal (HPL w/o group-DPO) induces a significantly more severe performance degradation than removing either the trajectory or step-level signals. This provides compelling evidence that the action-group granularity is the primary driver of HPL’s effectiveness, serving as a crucial bridge between coarse trajectory feedback and myopic action supervision and validating the synergistic benefit of our three-level approach.
@@ Appendix B Broader Impacts and Limitations
Our work on Hierarchical Preference Learning (HPL) presents a more efficient and effective paradigm for training capable autonomous agents from offline data. Technologically, this reduces the reliance on costly and often unsafe online exploration, making the development of sophisticated LLM agents more accessible and sustainable. The ability to learn from the compositional structure of tasks via action groups could lead to more reliable and predictable agents in real-world applications, from personalized assistants to complex workflow automation. On a societal level, the deployment of more competent agents can enhance productivity and assist with complex decision-making. However, as agent capabilities advance, ensuring their alignment with human values becomes paramount. The preferences learned from static datasets must be carefully curated to prevent the codification of biases or unintended behaviors. Furthermore, the increasing autonomy of such agents necessitates continued research into robust safety protocols, transparency, and ethical oversight to mitigate potential misuse and ensure these powerful technologies are developed responsibly for the benefit of society.
@@
Appendix C Experimental Details
C.1 Environments and Tasks
@@ C.1.3 InterCode-SQL
C.2 Models and Implementation Details
@@ C.3 Baselines
  • IPR (Xiong et al., 2024): A state-of-the-art process supervision method that performs step-level DPO using rewards estimated from Monte Carlo rollouts. IPR operates at the finest granularity, providing precise, localized feedback on individual actions. However, this myopic focus may overlook the synergistic value of multi-step sub-tasks.
C.4 Hyperparameters
@@
Appendix D Analysis
@@
Appendix E Additional Experiments
E.1 Sub-task Performance on ALFWorld
@@
E.2 Case Study
E.2.1 ALFWorld
@@ E.2.3 InterCode-SQL
Appendix F Prompts
F.1 ALFWorld

### v2

@@ 4.4 Ablation on Hierarchical DPO Losses (RQ4)
To investigate the individual contribution of each component in our hierarchical framework, we conduct an ablation study on the three DPO losses, with results shown in Figure 5. The results reveal that while all three loss components, trajectory, action, and group, contribute positively to the final performance, the group-level DPO is demonstrably the most critical. Across all benchmarks, removing the group-level signal (HPL w/o group-DPO) induces a significantly more severe performance degradation than removing either the trajectory or step-level signals. This provides compelling evidence that the action-group granularity is the primary driver of HPL’s effectiveness, serving as a crucial bridge between coarse trajectory feedback and action supervision and validating the synergistic benefit of our three-level approach.
@@ 5 Conclusion
Reproducibility Statement
Our experimental setup is briefly summarized in Section 4.1. To facilitate the reproduction of our results, Appendix D offers a detailed account of all the benchmarks, data generation process, baselines, and all hyperparameters. We also release the code and data, which are available at: https://github.com/PatrickG1014/HPL.
Acknowledgments
This work is supported in part by National Natural Science Foundation of China (No. 62472427 and No. 62422215), Major Innovation & Planning Interdisciplinary Platform for the “Double First-Class” Initiative, Renmin University of China, Public Computing Cloud, Renmin University of China, fund for building world-class universities (disciplines) of Renmin University of China, and in part by the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) under Grant No. GML-KF-24-33.
@@ Appendix B Broader Impacts and Limitations
Our work on Hierarchical Preference Learning (HPL) presents a more efficient and effective paradigm for training capable autonomous agents from fixed datasets after a one-shot exploration phase. Technologically, this reduces the reliance on costly and often unsafe online exploration, making the development of sophisticated LLM agents more accessible and sustainable. The ability to learn from the compositional structure of tasks via action groups could lead to more reliable and predictable agents in real-world applications, from personalized assistants to complex workflow automation. On a societal level, the deployment of more competent agents can enhance productivity and assist with complex decision-making. However, as agent capabilities advance, ensuring their alignment with human values becomes paramount. The preferences learned from static datasets must be carefully curated to prevent the codification of biases or unintended behaviors. Furthermore, the increasing autonomy of such agents necessitates continued research into robust safety protocols, transparency, and ethical oversight to mitigate potential misuse and ensure these powerful technologies are developed responsibly for the benefit of society.
@@
Appendix C Discussion
In this section, we discuss why group-level objectives are effective for long-horizon LLM agents. Building on both our theoretical analysis and empirical ablations, we highlight two complementary perspectives: (i) segmentation introduces a useful structural prior over sub-tasks, and (ii) group-level objectives have favorable statistical properties in the finite-data, limited-rollout regime.
Segmentation as a structural prior. Group-level supervision implicitly encodes prior knowledge about long-horizon task structure: instead of treating every individual action as an equally important learning unit, it focuses the objective on short sub-trajectories that are more likely to correspond to meaningful sub-tasks. Although HPL reuses exactly the same interaction data as trajectory- and step-level baselines, reorganizing this data into action groups changes where supervision is concentrated. Empirical results show that the semantic variant using a stronger segmentation prior achieves the best overall performance, while simpler segmenters still improve over both trajectory and step-level DPO, despite having almost no semantic information about sub-task boundaries.
Statistical properties in the finite-data regime. From an asymptotic perspective, a sufficiently expressive step-level DPO objective could in principle represent the same long-horizon credit assignment as a group-level objective. Our focus, however, is the practically relevant regime where both the dataset size and the MC rollout budget per state are limited. Under this regime, step-level MC estimates provide localized but noisy supervision: each decision point is updated from only a small number of rollouts. Group-level objectives alleviate this by aggregating multiple actions into a single supervised unit. With the same overall rollout budget, group-level rewards effectively average out noise over longer sub-trajectories and directly evaluate the joint contribution of several actions to task success. Proposition 1 formalizes this intuition through a bias–variance analysis, and our ablation studies further show that removing the group-level loss significantly degrades performance.
Appendix D Experimental Details
D.1 Environments and Tasks
@@ D.1.3 InterCode-SQL
D.2 Models and Implementation Details
@@ D.3 Baselines
  • IPR (Xiong et al., 2024): A state-of-the-art process supervision method that performs step-level DPO using rewards estimated from Monte Carlo rollouts. IPR operates at the finest granularity, providing precise, localized feedback on individual actions. However, this step-level focus may overlook the synergistic value of multi-step sub-tasks.
D.4 Hyperparameters
@@
D.5 Resource Comparison
We report in Table 5 a resource comparison of SFT, ETO, IPR, and the HPL variants on ALFWorld with Qwen2.5-1.5B-Instruct, including whether an external powerful LLM is used, the number of LLM calls, and the generation/training time.
Table 5: Resource comparison of SFT, ETO, IPR, and HPL variants on the ALFWorld benchmark with Qwen2.5-1.5B-Instruct.
Method External powerful LLM # LLM calls Gen time Train time
SFT 0 0 18min
ETO $\sim 30{,}000$ 1h 7min 13min
IPR $\sim 750{,}000$ (step-level part) 6h 13min (step-level part) 26min
HPL (Fixed-N(3)) $\sim 207{,}000$ (group-level part) 3h 35min (group-level part) 25min
HPL (Fixed-K(3)) $\sim 213{,}000$ (group-level part) 4h 15min (group-level part) 26min
HPL (Uncertainty) $\sim 194{,}000$ (group-level part) 3h 47min (group-level part) 28min
HPL (Semantic) $\sim 221{,}000$ (group-level part) 3h 21min (group-level part) 26min
During data generation, we adopt the same parallel sampling implementation provided by the ETO and IPR codebases to accelerate environment interaction. The actual generation time depends primarily on the degree of parallelism in environment rollouts, as well as the time required to reset the environment and progress it to the desired states; the LLM call latency is not the dominant factor. The reported training times are based on real runs using 4 NVIDIA A800 80G GPUs.
Appendix E Illustrative Example of Group Reward Estimation
In this section, we provide a concrete walkthrough of how the Monte Carlo (MC) rollout mechanism estimates the expected outcome reward for a specific action group, as defined in Equation 2.
Consider an example from the ALFWorld benchmark:
  • Task Instruction: put a clean apple in fridge
  • Current Context ($c_{i}$): The agent is located at the sinkbasin 1 and is holding a dirty apple 1.
  • Candidate Action Group ($G_{i}$): The policy generates a coherent sequence of actions intended to complete the “cleaning” sub-task:
    Gi = [go to sinkbasin 1, clean apple 1 with sinkbasin 1]
To estimate the reward $\hat{r}(G_{i})$, we first execute $G_{i}$. The simulation state transitions to a point where the agent is holding a clean apple at the sink. From this state, we perform $M=3$ stochastic rollouts using the reference policy $\pi_{ref}$ to see if the task can be successfully completed.
  • Rollout 1 ($\tau^{(1)}$): The agent successfully navigates to the fridge, opens it, and places the apple inside.
    $\rightarrow$ Outcome Reward $R(\tau^{(1)})=1.0$ (Success).
  • Rollout 2 ($\tau^{(2)}$): The agent navigates to the fridge but attempts to place the apple without opening the fridge first. It fails to recover within the step limit.
    $\rightarrow$ Outcome Reward $R(\tau^{(2)})=0.0$ (Failure).
  • Rollout 3 ($\tau^{(3)}$): The agent navigates to the fridge, opens it, and successfully places the apple.
    $\rightarrow$ Outcome Reward $R(\tau^{(3)})=1.0$ (Success).
Finally, the estimated reward for group $G_{i}$ is calculated as the average of these outcomes:
$\hat{r}(G_{i})=\frac{1}{3}(1.0+0.0+1.0)\approx 0.67.$
This value $\hat{r}(G_{i})=0.67$ serves as the quality label for this action group. If paired with a lower-quality group (e.g., one that failed to clean the apple), the difference $\Delta R$ determines the sample difficulty for our curriculum scheduler.
Appendix F Details of Dual-Layer Curriculum Scheduler
In this section, we provide the algorithmic implementation of our curriculum scheduler and a visual illustration of the phase-wise training progression.
Algorithm 1 outlines the logic for partitioning the preference data into the $3\times 3$ grid based on Group Length ($L$) and Sample Difficulty ($\Delta R$), and selecting the active data subsets for the current training phase $s$. We utilize the hyperparameters specified in Appendix D.4.
Algorithm 1 HPL Dual-Layer Curriculum Scheduler
1:Input: Dataset $\mathcal{D}_{all}$, Phase $s\in\{1,2,3\}$, Thresholds $T_{L},T_{\Delta R}$.
2:Output: Training subset $\mathcal{D}_{train}^{(s)}$.
3:Step 1: Data Partitioning
4: Partition $\mathcal{D}_{all}$ into $3\times 3$ buckets $\mathcal{B}_{l,d}$ where $l,d\in\{1,2,3\}$.
5: For each sample $x\in\mathcal{D}_{all}$, assign to $\mathcal{B}_{l,d}$ based on:
  • Length Level $l$: Determined by group length vs. $T_{L}$ (Short $\to$ 1, Long $\to$ 3).
  • Difficulty Level $d$: Determined by reward gap vs. $T_{\Delta R}$ (Easy $\to$ 1, Hard $\to$ 3).
6:Step 2: Phase-based Selection
7: Define the set of active bucket indices $\mathcal{I}^{(s)}$ for current phase $s$:
8:if $s=1$ then
9:  $\mathcal{I}^{(1)}\leftarrow\{(1,1)\}$ {Phase 1: Foundational (Short & Easy)}
10:else if $s=2$ then
11:  $\mathcal{I}^{(2)}\leftarrow\{(1,1),(1,2),(2,1)\}$ {Phase 2: Expansion}
12:else
13:  $\mathcal{I}^{(3)}\leftarrow\{(l,d)\mid 1\leq l,d\leq 3\}$ {Phase 3: Full Scale}
14:end if
15:return $\mathcal{D}_{train}^{(s)}\leftarrow\bigcup_{(l,d)\in\mathcal{I}^{(s)}}\mathcal{B}_{l,d}$
Appendix G Analysis
@@
Appendix H Additional Experiments
H.1 Sub-task Performance on ALFWorld
@@
H.2 Case Study
H.2.1 ALFWorld
@@ H.2.3 InterCode-SQL
Appendix I Prompts
I.1 ALFWorld